# A Bayesian Flow Network Framework for Chemistry Tasks

### v1

@@ 1 Introduction
Autoregressive models (ARs) including SMILES-based or fragment-based models1, 2, 3, 4 that leverage the power of language models (LMs) and graph-based models5, 6, 7 coupled with advanced techniques such as Monte Carlo tree search6, 7 have been proved their success in several de novo design benchmarks8, 4 consisted of drug-like molecules. The constraint of ARs, i.e., the number of sampling steps is the size of generated object, however, limits the potential of generating large molecules. Conversely, the recently emerging denoising-diffusion models9 (DMs) offer a way to generate objects of any size within a fixed sequence of sampling process. However, it has been pointed out in the research of C. Vignac et al10 that SMILES-based models generally worked better than graph DMs even when a dedicatedly designed discrete diffusion method was applied.
Bayesian flow networks11 (BFNs) are in a different category of generative models that decouple the sampling process with the size of generated objects as well. Different from DMs, BFNs directly work on the parameters of data distributions which naturally enable them to handle both continuous (including discretised) and discrete data without any data preprocessing or change of (mathematical) framework. Although the authors of BFN showed evidence in the origin paper11 that BFN advantaged over discrete DMs on discrete data generating, e.g., text generation, the recent researches considering de novo molecule design only successfully employed it on continuous and discretised data, e.g., 3D molecular conformation generation12 rather than language-like representations such as SMILES13 or SELFIES14. One potential reason discouraging the application to text generation is the lack of exact analytical expression for the accuracy schedule $\beta(t)$, one critical component of BFNs, in the discrete case, while the speculated quadratic $\beta(t)$ in the original paper is, as admitted by the authors11, suboptimal.
In this paper, we introduce ChemBFN, a Bayesian Flow Network framework for Chemistry tasks, that leverages our newly proposed accuracy schedule and transformer15 encoder model to generate 1D language-like molecular representations e.g., SMILES and SELFIES. The experiments demonstrated that models with our accuracy schedule outperform those with the quadratic accuracy schedule. Besides, the generative training of BFN method can be a powerful pretraining strategy for downstream tasks in molecular property predictions, including regressions and classifications, and reaction yield predictions.
@@ 2.1 Model Architecture
Our model is an adaptation of DiT16 model. The differences in our implementation include (1) the use of categorical distributions rather than image embeddings for input tokens because we are not dealing with images; (2) logits output that are then transformed into probabilities by softmax function; (3) the replacement of activation function with SELU17 function; (4) the use of a 2-layer MLP to form the time embedding since “time” in BFN is continuous from 0 to 1; (5) the employment of XPOS18 variation of rotary positional embedding19. The architecture is shown in Figure 1.
Figure 1: Visualised scheme of our model. The architecture is inspired by DiT16. The multi-head self-attention layers did not use causal masking which is the same as BERT20 while we replaced the commonly used positional embedding method (absolute positional embedding used in DiT, BERT and RoBERTa21 models) with the novel XPOS18 variation of rotary positional embedding19. Note that each FFN (feed-forward network) layer adsorbs a dropout layer.
Following the notations of the BFN paper11, the parameter of categorical distributions inputted into the neural network is denoted by $\boldsymbol{\theta}=(\theta^{(1)},\theta^{(2)},...,\theta^{(D)})\in[0,1]^{KD}$ ($K$ is the number of categories, $D$ is the number of input data, and $\theta^{(i)}$ is the $i^{th}$ data) and the output distribution at time step $t$ is denoted by $\boldsymbol{p}_{O}(\cdot|\boldsymbol{\theta};t)\in[0,1]^{KD}$. We denote the sum of time embedding vector and conditioning vector as $\boldsymbol{c}$. A null conditioning $\phi$ is equivalent to a zero vector $\boldsymbol{0}$.
In each experiment described in the later text, we employed the same hyperparameters of the model except category number $K$ that depends on molecular representations. The 2-layer MLP with SELU activation has the shape of [1, 256, 512]. We employed 12 Transformer layers, of which had 8 attention heads each, with the attention temperature22 $\tau=\sqrt{2d_{h}}$ ($d_{h}$ is the feature number of each attention head). The dropout rate was 0.01 and the hidden feature number was 512. These settings lead to a total learnable parameters of the model of the magnitude of 54M.
@@ 2.2 A New Accuracy Schedule
Ideally, an accuracy schedule function $\beta(t)$ drives the expectation of entropy of the input distribution $\mathbb{E}_{\boldsymbol{p}_{F}(\boldsymbol{\theta}|\boldsymbol{x};t)}H[\boldsymbol{p}_{I}(\boldsymbol{x}|\boldsymbol{\theta})]$ to decrease linearly with $t$, where $\boldsymbol{x}$ stands for the clear data, $\boldsymbol{p}_{F}(\boldsymbol{\theta}|\boldsymbol{x};t)$ represents Bayesian flow distribution, and $\boldsymbol{p}_{I}(\boldsymbol{x}|\boldsymbol{\theta})$ is the input distribution as denoted in the origin paper11. The mathematical difficulty of deriving the expectation analytically in the discrete case compels us to speculate from intuition. The authors of BFN claimed that “$\beta(t)=t^{2}\beta(1)$ was a reasonable approximation”, but disclosed later that finding a suitable value for the hyperparameter $\beta(1)$ was not an easy job11.
Here, we give our estimation of $\beta(t)$. If we estimate the expected entropy of the input distribution (denoted as $E$ for short) as $E\sim f(K)e^{-\frac{K}{4}\beta(t)}$, then the relationship $E(t)=(1-t)E(0)+tE(1)$ that eliminates the unknown function $f(K)$ gives us
@@ 2.3 Datasets and Benchmarks
Two benchmarks – MOSES8 and GuacaMol4 – were used to evaluate the generative performance, e.g., the similarity between generated molecules and training molecules, of ChemBFN. We reported the distribution-learning metrics of these benchmarks in Section 3.
@@
The QM923 dataset was employed to study the capability of conditional generation of our method. We randomly selected 110k molecules, before which 3054 invalid data were removed, with the triple $(\epsilon_{HOMO},\epsilon_{LUMO},\Delta\epsilon_{HOMO-LUMO})$ as the conditioning label to form the training set.
@@
In order to evaluate the downstream performance, 40M unique SMILES and 190M unique SMILES strings were randomly selected from ZINC1524 database that formed two pretraining sets. The model trained on the 40M set was finetuned on several regression (ESOL, FreeSolv, and Lipo) and classification (BBBP, BACE, ClinTox, and HIV) subsets of widely used MoleculeNet25 benchmark. Each dataset was split into training/validation/testing sets in the ratio of 80/10/10 following the scaffold splitting method proposed in DeepChem26 project. We reported ROC-AUC (area under receiver operating characteristic curve) for classification tasks and RMSE (root-mean squared error) for regression tasks in Section 3.
@@
The USPTO-50k27 dataset, Buchwald-Hartwig and Suzuki-Miyaura reaction yield datasets from high-throughput experiments (HTE) cleaned by P. Schwaller et al28 were employed to train the model to predict reaction yield. We report coefficient of determination (R2 score) on testing sets in Section 3.
@@
AqSolDB29, a more challenging solubility dataset containing more species than ESOL, was used to investigate the effect of the size of pretraining data. A training/validation/testing (80/10/10) split was generated using scaffold splitting method. Testing MAE (mean absolute error) and RMSE were reported in Section 3.
@@ 3.1 Unconditional Generation
Figure 4: Visualisation of the impact on training loss, reconstruction loss $L^{r}$ and cts time loss $L^{\infty}$ of different accuracy schedules with different values of $\beta(1)$. $L^{r}$ and $L^{\infty}$ were computed on 1k discretised steps after training.
@@
Table 1: Comparing scores of MOSES benchmark when varying $\beta(1)$ value of different accuracy schedules. $\uparrow$ indicates that the higher is better and $\downarrow$ stands for the contrary. The best results are in bold. We used a sampling step of 1k.
$\beta(1)$ Valid $\uparrow$ FCD $\downarrow$ SNN $\uparrow$ Frag $\uparrow$ Scaf $\uparrow$ Filters $\uparrow$ Novelty $\uparrow$
quadratic 0.15 0.893 3.438 0.559 0.985 0.095 0.982 0.900
0.0829 0.895 3.772 0.551 0.984 0.096 0.985 0.900
0.0448 0.899 3.902 0.561 0.988 0.089 0.986 0.887
ours 0.0829 0.900 2.731 0.563 0.990 0.091 0.987 0.886
0.0448 0.900 3.580 0.568 0.987 0.075 0.987 0.877
@@
Table 2: Scores of MOSES and GuacaMol benchmarks when different padding strategies were used during training. $\uparrow$ for higher is better and $\downarrow$ for contrary. The best results are in bold. We used a sampling step of 1k.
Strategy MOSES
Valid $\uparrow$ FCD $\downarrow$ SNN $\uparrow$ Frag $\uparrow$ Scaf $\uparrow$ Filters $\uparrow$ Novelty $\uparrow$
dynamic 0.900 2.731 0.563 0.990 0.091 0.987 0.886
global 0.916 2.730 0.565 0.990 0.094 0.987 0.880
GuacaMol
Valid $\uparrow$ Unique $\uparrow$ Novelty $\uparrow$ KL Divergence $\uparrow$ FCD $\uparrow$
dynamic 0.799 0.815 0.975 0.810 0.370
global 0.807 0.818 0.975 0.808 0.399
Finally, we trained models applying the above optimal settings (i.e., $\beta(1)=20.4054/K$ and global padding) on MOSES and GuacaMol datasets. Both SMILES and SELFIES versions were implemented. The comparison with published state-of-the-art (SOTA) models3, 1, 5, 2, 7, 10, 4 are summarised in Table 3, Table 4, and Table 5. We found that (1) except FCD, metrics of both SMILES version and SELFIES version were close to SOTA performance. (2) number of sampling step as expected affected the validity of generated molecules (for SMILES version only because SELFIES always gives valid molecules14), but dropping from 1k steps to 100 steps did not degrade the performance a lot. If lower validity is acceptable, only sampling 10 steps significantly reduce the computational time without much impact on other qualities. Larger FCD (in the term of GuacaMol is lower FCD score) is a hint that BFNs learn the grammar of molecules rather than the way of combining characters within the dataset.
Table 3: Testing metrics on MOSES test set compared with SOTA models. The metrics of all other models were copied from the original paper. $\uparrow$ for the higher is better. (10, 100, 1k) are the number of sampling steps. * for SELFIES version. The best results are in bold.
Method Valid $\uparrow$ Unique@1k $\uparrow$ Unique@10k $\uparrow$ IntDiv $\uparrow$ IntDiv2 $\uparrow$ Novelty $\uparrow$
ARs JTN-VAE3 1.0 1.0 1.0 0.855 0.850 0.913
LatentGAN1 0.897 1.0 0.997 0.857 0.851 0.950
GraphINVENT5 0.964 1.0 0.998 0.857 0.851
MolGPT2 0.994 1.0 0.857 0.851 0.797
DMs DiGress10 0.857 1.0 0.950
BFNs ChemBFN10 0.835 1.0 0.999 0.851 0.844 0.921
ChemBFN100 0.911 1.0 0.998 0.837 0.831 0.884
ChemBFN1k 0.916 1.0 0.998 0.836 0.830 0.880
@@
Table 4: Metrics on MOSES scaffold test set. Settings are the same as Table 3 while $\downarrow$ for the lower is better.
@@
ARs JTN-VAE3 0.938 0.519 0.995 0.101 0.976
LatentGAN1 0.828 0.513 0.997 0.107 0.974
GraphINVENT5 1.223 0.539 0.986 0.127 0.950
DMs DiGress10 1.19 0.52 0.148 0.971
BFNs ChemBFN10 2.768 0.533 0.988 0.145 0.976
ChemBFN100 2.604 0.562 0.991 0.103 0.985
ChemBFN1k 2.730 0.565 0.990 0.094 0.986
@@
Table 5: Testing metrics on GuacaMol distribution-learning tasks. Settings are the same as Table 3.
@@
ARs MolGPT2 0.981 0.998 1.0 0.992 0.907
SMILES LSTM4 0.959 1.0 0.912 0.991 0.913
VGAE-MCTS7 1.0 1.0 1.0 0.659 0.009
DMs DiGress10 0.852 1.0 0.999 0.929 0.680
BFNs ChemBFN1k 0.807 0.818 0.975 0.808 0.399

### v2

@@ Abstract
keywords:
molecule generation,molecular property prediction,reaction yield prediction
[…]
@@ 1 Introduction
Autoregressive models (ARs) including SMILES-based or fragment-based models1, 2, 3, 4, 5, 6, 7, 8, 9 that leverage the power of language models (LMs) and reinforcement learning7, 8, 9 and graph-based models10, 11, 12, 13, 14, 15 coupled with advanced techniques such as Monte Carlo tree search11, 12, 13 have been proved their success in several de novo design benchmarks16, 6 consisted of drug-like molecules. The constraint of ARs, i.e., the number of sampling steps is the size of generated object, however, limits the potential of generating large molecules. Conversely, the recently emerging denoising-diffusion models17 (DMs) offer a way to generate objects of any size within a fixed sequence of sampling process. However, it has been pointed out in the research of C. Vignac et al18 that SMILES-based models generally worked better than graph DMs even when a dedicatedly designed discrete diffusion method was applied.
Bayesian flow networks19 (BFNs) are in a different category of generative models that decouple the sampling process with the size of generated objects as well. Different from DMs, BFNs directly work on the parameters of data distributions which naturally enable them to handle both continuous (including discretised) and discrete data without any data preprocessing or change of (mathematical) framework. Although the authors of BFN showed evidence in the original paper19 that BFN advantaged over discrete DMs on discrete data generating, e.g., text generation, the recent researches considering de novo molecule design only successfully employed it on continuous and discretised data, e.g., 3D molecular conformation generation20 rather than language-like representations such as SMILES21 or SELFIES22. One potential reason discouraging the application to text generation is the lack of exact analytical expression for the accuracy schedule $\beta(t)$, one critical component of BFNs, in the discrete case, while the speculated quadratic $\beta(t)$ in the original paper is, as admitted by the authors19, suboptimal.
In this paper, we introduce ChemBFN, a Bayesian Flow Network framework for Chemistry tasks, that leverages our newly proposed accuracy schedule and transformer23 encoder model to generate 1D language-like molecular representations e.g., SMILES and SELFIES. The experiments demonstrated that models with our accuracy schedule outperform those with the quadratic accuracy schedule. Besides, the generative training of BFN method can be a powerful pretraining strategy for downstream tasks in molecular property predictions, including regressions and classifications, and reaction yield predictions.
@@ 2 Methods
2.1 Discrete Bayesian Flow Networks
A functional BFN is consisted of a neural network (NN) model that converts the input distribution $\boldsymbol{p}_{I}(\boldsymbol{x}|\boldsymbol{\theta})$ into the output distribution $\boldsymbol{p}_{O}(\boldsymbol{x}|\boldsymbol{\theta};t)$ and a Bayesian update process that updates the pervious input distribution to the current state according to a sender distribution $\boldsymbol{p}_{S}(\boldsymbol{y}|\boldsymbol{x};\alpha)$, where $\boldsymbol{\theta}$ is the parameter of data $\boldsymbol{x}$, and $\boldsymbol{y}$ is a sample of $\boldsymbol{x}$19. The none-negative monotonic increasing function $\alpha$, namely accuracy rate, guides the sender distribution to moving to a more informative direction along with the time19. Since $\alpha$ can be either continuous or discretised, a continuous accuracy schedule $\beta(t)$ is defined instead, which generates $\alpha$ as
$\alpha=\left\{\begin{array}[]{ll}{\dfrac{d}{dt}\beta(t)},&{\textrm{when $\alpha$ is continuous}}\\ {\beta(t_{i})-\beta(t_{i-1})},&{\textrm{when $\alpha$ is discretised}}.\end{array}\right.$ (1)
In the discrete case, we have (1) all the distributions are K-class categorical distributions; (2) the sample is defined as $\boldsymbol{y}=\mathcal{N}(\alpha(K\boldsymbol{e_{x}}-\boldsymbol{1}),\alpha K\boldsymbol{I})$ when a Gaussian sampling is utilised, where $\boldsymbol{e_{x}}$ is the one-hot representation of data $\boldsymbol{x}$; (3) the Bayesian update function is defined as $h(\theta^{(d)},y^{(d)},\alpha)=e^{y^{(d)}}\theta^{(d)}/\sum_{k=1}^{K}e^{y^{(d)}_{k}}\theta^{(d)}_{k}$, where $\cdot^{(d)}$ is the $d^{th}$ parameter19.
During the training stage, a receiver distribution $\boldsymbol{p}_{R}(\hat{\boldsymbol{y}}|\boldsymbol{\theta};t,\alpha)$ is drawn by sampling the output of NN with the same sampling method as sender distribution19. The model is optimised by minimising the Kullback-Leibler divergence between the receiver distribution and the sender distribution, which is decoupled as a $n$-step loss ($L^{n}$) and a reconstruction loss ($L^{r}$) and only the first loss in practice is used19. The limit case, i.e, continuous time loss, $L^{\infty}=\lim_{n\rightarrow\infty}L^{n}$ has been proved more efficient19. During the sampling (generating) stage, since the receiver distribution has been trained to match the sender distribution, i.e., $\boldsymbol{p}_{R}(\hat{\boldsymbol{y}}|\boldsymbol{\theta};t,\alpha)\sim\boldsymbol{p}_{S}(\boldsymbol{y}|\boldsymbol{x};\alpha)$, the receiver distribution is used in Bayesian update process directly to update the input distribution (initialised as a uniform distribution over K categories) where a discretised $\alpha$ is employed.
@@ 2.2 Model Architecture
Our model is an adaptation of DiT24 model. The differences in our implementation include (1) the use of categorical distributions rather than image embeddings for input tokens because we are not dealing with images; (2) logits output that are then transformed into probabilities by softmax function; (3) the replacement of activation function with SELU25 function; (4) the use of a 2-layer multilayer perceptron (MLP) to form the time embedding since “time” in BFN is continuous from 0 to 1; (5) the employment of XPOS26 variation of rotary positional embedding27. The architecture is shown in Figure 1.
Figure 1: Visualised scheme of our model. The architecture is inspired by DiT24. The multi-head self-attention layers did not use causal masking which is the same as BERT28 while we replaced the commonly used positional embedding method (absolute positional embedding used in DiT, BERT and RoBERTa29 models) with the novel XPOS26 variation of rotary positional embedding27. Note that each FFN (feed-forward network) layer adsorbs a dropout layer.
Following the notations of the BFN paper19, the parameter of categorical distributions inputted into the neural network is denoted by $\boldsymbol{\theta}=(\theta^{(1)},\theta^{(2)},...,\theta^{(D)})\in[0,1]^{KD}$ ($K$ is the number of categories, $D$ is the number of input data, and $\theta^{(d)}$ is the $d^{th}$ parameter) and the output distribution at time step $t$ is denoted by $\boldsymbol{p}_{O}(\cdot|\boldsymbol{\theta};t)\in[0,1]^{KD}$. We denote the sum of time embedding vector and conditioning vector as $\boldsymbol{c}$. A null conditioning $\phi$ is equivalent to a zero vector $\boldsymbol{0}$.
In each experiment described in the later text, we employed the same hyperparameters of the model except category number $K$ that depends on molecular representations. The 2-layer MLP with SELU activation has the shape of [1, 256, 512]. We employed 12 Transformer layers, of which had 8 attention heads each, with the attention temperature $\tau=\sqrt{2d_{h}}$ ($d_{h}$ is the feature number of each attention head)30. The dropout rate was 0.01 and the hidden feature number was 512. These settings lead to a total learnable parameters of the model of the magnitude of 54M.
@@ 2.3 A New Accuracy Schedule
In the case of BFN, an accuracy schedule function $\beta(t)$ drives the expectation of entropy of the input distribution $\mathbb{E}_{\boldsymbol{p}_{F}(\boldsymbol{\theta}|\boldsymbol{x};t)}H[\boldsymbol{p}_{I}(\boldsymbol{x}|\boldsymbol{\theta})]$ to decrease linearly with $t$, where $\boldsymbol{x}$ stands for the clear data, $\boldsymbol{p}_{F}(\boldsymbol{\theta}|\boldsymbol{x};t)$ represents Bayesian flow distribution, and $\boldsymbol{p}_{I}(\boldsymbol{x}|\boldsymbol{\theta})$ is the input distribution as denoted in the original paper19. The mathematical difficulty of deriving the expectation analytically in the discrete case compels us to speculate from intuition. The authors of BFN claimed that “$\beta(t)=t^{2}\beta(1)$ was a reasonable approximation”, but disclosed later that finding a suitable value for the hyperparameter $\beta(1)$ was not an easy job19.
Here, we give our estimation of $\beta(t)$. If we estimate the expected entropy of the input distribution (denoted as $E$ for short) as $E\sim f(K)e^{-\frac{K}{4}\beta(t)}$, then the relationship $E(t)=(1-t)E(0)+tE(1)$ that eliminates the unknown factor $f(K)$ gives us
@@ 2.4 Datasets and Benchmarks
Two benchmarks – MOSES16 and GuacaMol6 – were used to evaluate the generative performance, e.g., the similarity between generated molecules and training molecules, of ChemBFN. We reported the distribution-learning metrics of these benchmarks in Experiments and Results. A summary of these metrics is in Table 1.
Table 1: A brief summary of used metrics of MOSES and GuacaMol benchmarks
Metrics Description
Valid Fraction of valid molecules
Unique Fraction of unique molecules
IntDiv1 and IntDiv2 Internal diversities
Novelty Fraction of generated unseen molecules compared with training data
FCD Fréchet ChemNet Distance31
SNN Tanimoto similarity to a nearest neighbour
Frag BRICS fragment32 cosine similarity
Scaf Bemis–Murcko scaffold33 cosine similarity
Filter Fraction of molecules that fit pre-defined constructions
KL Divergence Kullback-Leibler divergence
@@
The QM934 dataset was employed to study the capability of conditional generation of our method. We randomly selected 110k molecules, before which 3054 invalid data were removed, with the triple $(\epsilon_{HOMO},\epsilon_{LUMO},\Delta\epsilon_{HOMO-LUMO})$ as the conditioning label to form the training set.
@@
In order to evaluate the downstream performance, 40M unique SMILES and 190M unique SMILES strings were randomly selected from easily accessed ZINC1535 database that formed two pretraining sets. The model trained on the 40M set was finetuned on several regression (ESOL, FreeSolv, Lipo, etc.) and classification (BBBP, BACE, ClinTox, HIV, etc.) tasks, including the subsets of widely used MoleculeNet36 benchmark. A brief description of used MoleculeNet tasks is in Table 2. Each dataset was split into training/validation/testing sets in the ratio of 80/10/10 following the scaffold splitting method proposed in DeepChem37 project. We reported ROC-AUC (area under receiver operating characteristic curve) for classification tasks and RMSE (root-mean squared error) for regression tasks in Experiments and Results. In addition to the tasks of MoleculeNet, two less biased datasets – the public ADME dataset published by C. Fang et al38 consisted of 6 dedicatedly collected absorption, distribution, metabolism, and excretion (ADME) in vitro endpoints together with a Kinase inhibitor dataset prepared by J.Wu et al39 that contains bioactivities of total 141,086 compounds for 354 kinases – were employed to further benchmark our method in activity prediction. A brief summary of sub-tasks of ADME dataset is in Table 2. For ADME dataset We employed the same split provided by C. Fang et al38; for Kinase inhibitor dataset, we prepared a random split and a scaffold split (both had training/validation/testing = 80/10/10). The testing MAE, RMSE, Pearson’s correlation coefficient (R value), and averaged ROC-AUC were reported in Experiments and Results.
Table 2: A brief summary of used MoleculeNet and public ADME tasks
Name № molecules № tasks Label
ESOL 1,128 1 Aqueous solubility $log_{10}(S/{\rm mol\cdot L^{-1}})$
FreeSolv 642 1 Experimental hydration free energy / kcal/mol
Lipo 4,200 1 Octanol/water distribution coefficient $log_{10}D_{7.4}$
HLM 3,087 1 Logarithm of human liver microsomal stability / $\rm mL\cdot min^{-1}\cdot kg^{-1}$
RLM 3,054 1 Logarithm of rat liver microsomal stability / $\rm mL\cdot min^{-1}\cdot kg^{-1}$
hPPB 1,808 1 Logarithm of human plasma protein binding (percent unbound)
rPPB 884 1 Logarithm of rat plasma protein binding (percent unbound)
[…]
Solubility 2,173 1 Aqueous solubility $log_{10}(S/{\rm\mu g\cdot mL^{-1}})$ at PH = 6.8
BBBP 2,039 1 If a compound penetrates the blood-brain barrier
BACE 1,513 1 If a compound inhibits BACE-1 protein
ClinTox 1,478 2 Task 1: if a drug has been approved by FDA
Task 2: if the same drug had toxicity during clinical trail
HIV 41,127 1 If a compound is an HIV inhibitor
@@
The USPTO-50k40 dataset, Buchwald-Hartwig and Suzuki-Miyaura reaction yield datasets from high-throughput experiments (HTE) cleaned by P. Schwaller et al41 were employed to train the model to predict reaction yield. USPTO-50k that contains 50k reactions mined from patents were used to pre-train the model while HTE data were used for fine-tuning. We report coefficient of determination (R2 score) on testing sets in Experiments and Results.
@@
AqSolDB42, a more challenging solubility dataset containing more species than ESOL, was used to investigate the effect of the size of pretraining data. A training/validation/testing (80/10/10) split was generated using scaffold splitting method. Testing MAE (mean absolute error) and RMSE were reported in Experiments and Results.
@@ 3.1 Unconditional Generation
Figure 4: Visualisation of the impact on training loss, reconstruction loss $L^{r}$ and continuous (cts) time loss $L^{\infty}$ of different accuracy schedules with different values of $\beta(1)$. $L^{r}$ and $L^{\infty}$ were computed on 1k discretised steps after training.
@@
Table 3: Comparing scores of MOSES benchmark when varying $\beta(1)$ value of different accuracy schedulesa
[…]
[][c] $\beta(1)$ Valid $\uparrow$ FCD $\downarrow$ SNN $\uparrow$ Frag $\uparrow$ Scaf $\uparrow$ Filters $\uparrow$ Novelty $\uparrow$ quad 0.15 0.893 $\pm$ 0.001 3.438 $\pm$ 0.034 0.559 $\pm$ 0.000 0.985 $\pm$ 0.000 0.095 $\pm$ 0.001 0.982 $\pm$ 0.000 0.900 $\pm$ 0.002 0.0829 0.895 $\pm$ 0.001 3.772 $\pm$ 0.012 0.551 $\pm$ 0.001 0.984 $\pm$ 0.001 0.096 $\pm$ 0.006 0.985 $\pm$ 0.001 0.900 $\pm$ 0.002 0.0448 0.899 $\pm$ 0.003 3.902 $\pm$ 0.045 0.561 $\pm$ 0.000 0.988 $\pm$ 0.000 0.089 $\pm$ 0.006 0.986 $\pm$ 0.001 0.887 $\pm$ 0.003 ours 0.0829 0.900 $\pm$ 0.001 2.731 $\pm$ 0.015 0.563 $\pm$ 0.000 0.990 $\pm$ 0.000 0.091 $\pm$ 0.004 0.987 $\pm$ 0.001 0.886 $\pm$ 0.000 0.0448 0.900 $\pm$ 0.001 3.580 $\pm$ 0.008 0.568 $\pm$ 0.000 0.987 $\pm$ 0.000 0.075 $\pm$ 0.006 0.987 $\pm$ 0.000 0.877 $\pm$ 0.001
  • a
    $\uparrow$ indicates that the higher is better and $\downarrow$ stands for the contrary. The best results are in bold. We used a sampling step of 1k.
@@
Table 4: Scores of MOSES and GuacaMol benchmarks when different padding strategies were used during traininga
[…]
[][c] Strategy MOSES Valid $\uparrow$ FCD $\downarrow$ SNN $\uparrow$ Frag $\uparrow$ Scaf $\uparrow$ Filters $\uparrow$ Novelty $\uparrow$ dynamic 0.900 $\pm$ 0.001 2.731 $\pm$ 0.015 0.563 $\pm$ 0.000 0.990 $\pm$ 0.000 0.091 $\pm$ 0.004 0.987 $\pm$ 0.001 0.886 $\pm$ 0.000 global 0.916 $\pm$ 0.001 2.730 $\pm$ 0.014 0.565 $\pm$ 0.001 0.990 $\pm$ 0.000 0.094 $\pm$ 0.002 0.987 $\pm$ 0.001 0.880 $\pm$ 0.002 GuacaMol Valid $\uparrow$ Unique $\uparrow$ Novelty $\uparrow$ KL Divergence $\uparrow$ FCD $\uparrow$ dynamic 0.799 $\pm$ 0.003 0.815 $\pm$ 0.002 0.975 $\pm$ 0.000 0.810 $\pm$ 0.001 0.370 $\pm$ 0.003 global 0.807 $\pm$ 0.003 0.818 $\pm$ 0.001 0.975 $\pm$ 0.001 0.808 $\pm$ 0.010 0.399 $\pm$ 0.002
  • a
    $\uparrow$ for higher is better and $\downarrow$ for contrary. The best results are in bold. We used a sampling step of 1k.
Finally, we trained models applying the above optimal settings (i.e., $\beta(1)=20.4054/K$ and global padding) on MOSES and GuacaMol datasets. Both SMILES and SELFIES versions were implemented. The comparison with published state-of-the-art (SOTA) models5, 3, 10, 4, 12, 18, 6 are summarised in Table 5, Table 6, and Table 7. We found that (1) except FCD, metrics of both SMILES version and SELFIES version were close to SOTA performance. (2) number of sampling step as expected affected the validity of generated molecules (for SMILES version only because SELFIES always gives valid molecules22), but dropping from 1k steps to 100 steps did not degrade the performance a lot. If lower validity is acceptable, only sampling 10 steps significantly reduce the computational time without much impact on other qualities. Larger FCD (in the term of GuacaMol is lower FCD score where FCD score = $e^{-0.2\rm FCD}$) is a hint that BFNs learn the grammar of molecules rather than the way of combining characters within the dataset.
Table 5: Testing metrics on MOSES test set compared with SOTA modelsa
Method Valid $\uparrow$ Unique@1k $\uparrow$ Unique@10k $\uparrow$ IntDiv1 $\uparrow$ IntDiv2 $\uparrow$ Novelty $\uparrow$
ARs JTN-VAE5 1.0 $\pm$ 0.0 1.0 $\pm$ 0.0 1.0 $\pm$ 0.0 0.855 $\pm$ 0.003 0.850 $\pm$ 0.004 0.913 $\pm$ 0.006
LatentGAN3 0.897 $\pm$ 0.003 1.0 $\pm$ 0.0 0.997 $\pm$ 0.000 0.857 $\pm$ 0.001 0.851 $\pm$ 0.001 0.950 $\pm$ 0.001
GraphINVENT10 0.964 1.0 0.998 0.857 0.851
MolGPT4 0.994 1.0 0.857 0.851 0.797
DMs DiGress18 0.857 1.0 0.950
BFNs ChemBFN10 0.835 $\pm$ 0.003 1.0 $\pm$ 0.0 0.999 $\pm$ 0.000 0.851 $\pm$ 0.000 0.844 $\pm$ 0.000 0.921 $\pm$ 0.002
ChemBFN100 0.911 $\pm$ 0.002 1.0 $\pm$ 0.0 0.998 $\pm$ 0.000 0.837 $\pm$ 0.000 0.831 $\pm$ 0.000 0.884 $\pm$ 0.002
ChemBFN1k 0.916 $\pm$ 0.001 1.0 $\pm$ 0.0 0.998 $\pm$ 0.000 0.836 $\pm$ 0.000 0.830 $\pm$ 0.000 0.880 $\pm$ 0.002
@@
  • a
    The metrics of all other models were copied from the original paper. $\uparrow$ for the higher is better. (10, 100, 1k) are the number of sampling steps. * for SELFIES version. The best results are in bold.
Table 6: Metrics on MOSES scaffold test seta
@@
ARs JTN-VAE5 0.938 $\pm$ 0.053 0.519 $\pm$ 0.007 0.995 $\pm$ 0.000 0.101 $\pm$ 0.011 0.976 $\pm$ 0.002
LatentGAN3 0.828 $\pm$ 0.012 0.513 $\pm$ 0.000 0.997 $\pm$ 0.001 0.107 $\pm$ 0.010 0.974 $\pm$ 0.001
GraphINVENT10 1.223 0.539 0.986 0.127 0.950
DMs DiGress18 1.19 0.52 0.148 0.971
BFNs ChemBFN10 2.768 $\pm$ 0.035 0.533 $\pm$ 0.000 0.988 $\pm$ 0.000 0.145 $\pm$ 0.004 0.976 $\pm$ 0.001
ChemBFN100 2.604 $\pm$ 0.040 0.562 $\pm$ 0.001 0.991 $\pm$ 0.000 0.103 $\pm$ 0.005 0.985 $\pm$ 0.001
ChemBFN1k 2.730 $\pm$ 0.014 0.565 $\pm$ 0.001 0.990 $\pm$ 0.000 0.094 $\pm$ 0.002 0.987 $\pm$ 0.001
@@
  • a
    Settings are the same as Table 5 while $\downarrow$ for the lower is better.
Table 7: Testing metrics on GuacaMol distribution-learning tasksa
@@
ARs MolGPT4 0.981 0.998 1.0 0.992 0.907
SMILES LSTM6 0.959 1.0 0.912 0.991 0.913
VGAE-MCTS12 1.0 1.0 1.0 0.659 0.009
DMs DiGress18 0.852 1.0 0.999 0.929 0.680
BFNs ChemBFN1k 0.807 $\pm$ 0.003 0.818 $\pm$ 0.001 0.975 $\pm$ 0.001 0.808 $\pm$ 0.010 0.399 $\pm$ 0.002